# Conformal four-point ladder integrals in diverse dimensions and polylogarithms

### v1

@@ Abstract
In the paper, the family of conformal four-point ladder diagrams in arbitrary space-time dimensions is considered. We use the representation obtained via explicit calculation using the operator approach and conformal quantum mechanics to study their properties, such as symmetries, loop and dimensional shift identities. In even integer dimensions, latter allows one to reduce the problem to two-dimensional case, where the notable factorization holds. Additionally, for a specific choice of propagator powers, we show that the representation can be written in the form of linear combinations of classical polylogarithms (with coefficients that are rational functions) and explore the structure of the resulting expressions.
Contents
Introduction
@@
In recent years, significant progress has been made in the computation of scattering amplitudes in various quantum field theories (see e. g. [1, 2, 3, 4, 5, 6]). In calculation of physical observables, e.g. in QCD, the algoritmizability and uniformity are often prioritized over the preserving specific symmetries at the intermediate steps. As a result, such techniques as integration-by-parts (IBP) [7], or differential equation methods [8, 9, 10], received significant development. On the other hand, in theories with enhanced symmetries, it is sometimes possible to exploit these symmetries directly in computations. A well-known example of such a theory is $\mathcal{N}=4$ SYM, especially in the planar limit. Underlying superconformal symmetry made it possible to fully constrain tree-level MHV amplitudes in a compact form [11, 12] even in the early years of study.
@@
Remarkably, the idea of using specific symmetries in Feynman integrals was implemented even before the establishment of IBP-reduction (which is nowadays widely used as a universal tool), leading to an equally universal approach. Special mention deserves a Gegenbauer polynomial technique [13, 14] (see also [15] for a modern review). It utilizes the $D$-dimensional rotational symmetry, which is the signature of all Feynman integrals. Despite challenges in its algorithmic implementation, this technique provided a possibility to obtain a series of important results [16, 17, 18] and continues to reveal deep connections with the other methods [19].
@@
However, not all symmetries of the theory can be read directly from its Lagrangian formulation. A notable example is the conjectured $\text{AdS}/\text{CFT}$ correspondence, which relates planar $\mathcal{N}=4$ SYM to string theory on $\text{AdS}_{5}\times S^{5}$. This duality in turn reveals an underlying integrable structure. For instance, in the so-called fishnet theory, which can be obtained as a deformation of planar $\mathcal{N}=4$ SYM [20, 21, 22, 23, 24], underlying integrability leads to significant results in the calculation of four-point Basso-Dixon correlators [25, 26, 27, 28, 29, 30, 31, 32, 33] and more general correlators [34, 35, 36, 37, 38, 39, 40]. Another hidden symmetry, crucial to the structure of planar $\mathcal{N}=4$ SYM, is the dual conformal symmetry [41, 42, 43], which together with ordinary conformal symmetry forms a Yangian symmetry [44]. This structure not only constrains tree-level amplitudes but can also be extended to the loop level, as reflected in the BDS ansatz for all-loop MHV amplitudes [45]. Note that dual conformal symmetry also finds its application beyond the four-dimensional theories [46, 47, 48, 49].
@@
From a technical point of view, dual conformal symmetry manifests as conformal invariance of the integrals associated with the dual Feynman graphs. These integrals have attracted attention since the early days of multiloop calculations and continue to be actively studied [50, 51]. Conformal symmetry at the level of perturbative integrals make it possible to use the powerful analytical identities, such as e.g. the star–triangle relation [52, 53, 54] or theory of graphical functions [55, 56], to simplify calculations dramatically. The star–triangle relation (interpreted as the Yang-Baxter equation; for a brief review see the subsection 5.3 in [57]) allows one to consider not only individual diagrams but entire families, such as fishnet diagrams [52], or conformal ladder diagrams in arbitrary dimensions $D$ [58, 59]. Recall that the ladder diagrams first studied by Ussyukina and Davydychev, who derived all-loop results in four dimensions [60, 61]. Later, a generalization of conformal ladder integrals were analyzed, and their analytic properties, such as single-valuedness, were explored in [62].
@@
Note that in the paper [54] we found the general $D$-dimensional conformal invariant solution of the Yang-Baxter equation, which generalizes $1$-dimensional solution of [59] and $2$-dimensional solution of [63], and underlies Lipatov’s integrable models describing high-energy behaviour of QCD as well as integrable structures of scattering amplitudes in $\mathcal{N}=4$ SUSY [64, 65, 66]. This Yang-Baxter solution was used, for example, in the investigations of the Yangian symmetries of perturbative integral in fish-net type, loom and checkerboard CFT [67, 68, 69].
@@
In one of our previous works [70], we used the graph-building operator technique [20, 21, 22, 27, 71, 28, 29, 72], together with the connection to conformal quantum mechanics [58, 59], to obtain an all-loop result for the conformal ladder and zig-zag four-point correlators in arbitrary dimensions. While our expression for ladder diagrams was fully analytical and valid in any dimension, it was formulated in terms of Gegenbauer polynomials, so despite its generality the form of the answer could be subtle for the practical use. Recently, remarkable progress has been made in two dimensions, where conformal ladder integrals were shown to be related to twisted partition functions [73, 74, 75], allowing them to be rewritten in terms of classical polylogarithms. In the present work, we show that in arbitrary even dimensions and specific choice of propagator indices our previous representation can be systematically expressed using classical polylogarithms and rational functions. Furthermore, we verify that our representation satisfies loop and dimensional shift identity, studied in [76, 74, 75]. The latter, among other things, allows one to express the answer for $D$-dimensional diagram in terms of two-dimensional, where the notable factorization holds (see [27, 19] for examples of such a factorization). We believe that these results can be useful not only from the practical point of view but also in revealing possible underlying symmetries, such as antipodal self-duality [77] in the fishnet theory which also holds for the one-loop ladder diagrams in $D=4$. Before closing this section we would like to mention that despite the rich variety of various results the studies of conformal integrals continue to attract interest [78, 79, 80].
@@
The paper is organized as follows. In Section 2 we review the calculation of the general family of conformal four-point ladder diagrams using graph-building operator and conformal quantum mechanics. After discussing the most general choice of parameters for the ladder diagrams we restrict ourselves to the case, described by a single parameter $\beta$. We show that obtained representation admits a dimensional shift $D\mapsto D+2$ for the general choice of $\beta$. Additionally, we show that the combination of dimensional shift operator with the graph-building operator allows us to construct the operator which increases number of loops $L\mapsto L+1$. In Section 3 we move to the case of ladder diagrams with specific parameter $\beta=1$ and present the derivation of the answer, containing only classical polylogarithms and rational functions. Then, we study representation based on two-dimensional factorized form, in the case $\beta=2,3,\ldots$. The last section contains our conclusions. Appendices are dedicated to discussion of technical details, namely in Appendix A we derive the answer for two-dimensional conformal ladder integrals using representation with Gegenbauer polynomials, in Appendix B we give the explicit check of the loop shift identity for four-dimensional diagram with $\beta=1$, Appendix C shows the properties of the answer in arbitrary dimensional case with $\beta=1$ and in Appendix D we discuss the details of two-dimensional factorization.
Conformal four-point ladder integrals in diverse dimensions
Formulation of the problem and introductory remarks
@@
To fix notation, in this section we recall some known facts (see e.g. Refs. [58, 70]). For massless $\varphi^{3}$ theory, we consider D-dimensional $L$-loop ladder integrals with arbitrary indices on the lines $\alpha_{k},\beta_{k},\gamma_{k}$:
@@
These integrals correspond to the momentum-space Feynman diagrams depicted in Fig. 1 (here the integrations are performed over the loop momenta $p_{i}$ (i=1,,Li=1,...,L)). The diagram given in Fig. 1 is presented in the dual form in Fig. 2. Here the integrations are performed over boldface vertices which are placed in the boxes of the diagram in Fig. 1.
@@
Figure 2: Dual graphical representation of the integral (2.1) written in the form (2.2). The integrations are performed over boldface vertices and each line with index $\alpha$: $x_{i}\stackrel{{\scriptstyle\alpha}}{{-\!\!\!-\!\!\!-\!\!\!-}}x_{j}$, corresponds to the propagator $x_{ij}^{-2\alpha}$
@@
For further applications, we fix the indices in the integral (2.2), which is depicted in Fig.2, by the conformal conditions for all boldface vertices [58, 70]
@@
When $\alpha_{k},\beta_{k},\gamma_{k}$ are positive real numbers the condition (2.3) ensures the convergence of the integral (2.2). Under the conformal transformations $y_{\mu}\to\frac{y_{\mu}}{y^{2}}$ (inversion of vector $y\in\mathbb{R}^{D}$; $\mu=1,...,D$ is the space index) we have for vectors $x_{i},x_{j}\in\mathbb{R}^{D}$:
@@
and in the case when the conditions (2.3) are fulfilled the integral (2.2) transforms as follows
@@
Thus, the function [^1] We use notation $(x_{ij})^{A}=(x_{ij}^{2})^{A/2}$.
@@
and $f_{(0)}=1$. Further, for simplicity we fix the parameters $\beta_{k}=D/2-\gamma_{k}\equiv\beta$ (which gives $\alpha_{k}=\beta$). In this case formula (2.6) simplifies to
@@
and these functions (associated with the diagrams in Fig 2) give contributions to the $4$-point amplitude in general $D$-dimensional bi-scalar fishnet CFT proposed in [22, 68]. For this choice of the parameters, we found in [58, 59, 70] that the generating function of $L$-loop ladder integrals (2.2) (with special normalization (2.7)) is represented in the form of a Green’s function for conformal quantum mechanics [^2] Here, to avoid confusion with the standard notation of cross-ratios, we use the notation $\sf x$} and $\sf y$} instead of $u$ and $w$ in [70] and $u$ and $v$ in [58].
$\begin{array}[]{c}\displaystyle\frac{a(\beta)}{({\sf x}-{\sf y})^{2(D/2-\beta)}}\cdot\sum_{L=0}^{\infty}\bigl(g\,a(\beta)\bigr)^{L}\,f_{(L)}(u,v;\beta)=\bigl\langle{\sf x}\bigr|\,\frac{1}{\hat{p}^{2\beta}-g\,\hat{q}^{-2\beta}}\,\bigl|{\sf y}\bigr\rangle\;,\\[5.69046pt] \hat{p}^{2}=\hat{p}_{\mu}\hat{p}_{\mu}\,,\;\;\;\hat{q}^{2}=\hat{q}_{\mu}\hat{q}_{\mu}\,,\;\;\;[\hat{q}_{\mu},\;\hat{p}_{\nu}]=i\delta_{\mu\nu}\,,\;\;\;\;{\sf x}:=\frac{1}{x_{12}}-\frac{1}{x_{42}}\,,\;\;\;\;{\sf y}:=\frac{1}{x_{32}}-\frac{1}{x_{42}}\,,\\[8.5359pt] \dfrac{a(\beta)}{(x-y)^{2(D/2-\beta)}}=\langle x|\dfrac{1}{\hat{p}^{2\beta}}|y\rangle\,\qquad(\forall x,y\in\mathbb{R}^{D}),\;\;\;\;\;\;a(\beta):=\dfrac{\Gamma(D/2-\beta)}{2^{2\beta}\,\pi^{D/2}\Gamma(\beta)}\;,\end{array}$ (2.8)
where $\{\hat{q}_{\mu},\;\hat{p}_{\nu}\}$ – are generators of $D$-dimensional Heisenberg algebra, and [70]
@@
Here, ${\sf x}^{\mu_{1}...\mu_{n}}$ is the traceless symmetric tensor with components that are homogeneous in ${\sf x}^{\mu}$ polynomials (see Appendix A in [70]); functions
@@
were introduced in [70] as the weights in the completeness condition of the eigenfunctions of operators $H_{\beta}=\hat{p}^{2\beta}\hat{q}^{2\beta}$; operators $H_{\beta}$ form a commutative set for all $\beta$ [58] while $\tau_{n,\nu}(\beta;\lambda)$ are eigenvalues of $H_{\beta}$ (for further details see subsection 2.3):
@@
Note also that we have relations between variables ${\sf x},{\sf y}$ (introduced in (2.8)) and cross-ratios (2.5)
@@
The representation (2.8) of the generating function of $L$-loop ladder integrals as a Green’s function for conformal quantum mechanics is very useful since we, for example, immediately deduce the symmetry properties of $\Phi^{(\beta)}_{L}({\sf x},{\sf y})$ [58]:
@@
which also follows from (2.9). Comparing (2.9) and (2.15) we obtain [70]
@@
Remark 1. Under the translation of variable $\nu\mapsto\nu-{\sf i}\frac{\beta}{2}$, the function (2.18) transforms to the expression
$\Phi^{(\beta)}_{L}({\sf x},{\sf y})=$ $\,(L!\,4^{L})4^{-\beta(L+1)}\sum\limits_{n=0}^{\infty}\mu(n)\,\frac{{\sf x}^{\mu_{1}...\mu_{n}}\,{\sf y}^{\mu_{1}...\mu_{n}}}{({\sf x}^{2}\,{\sf y}^{2})^{(D/4+n/2)}}\qquad\times$
$\qquad\times\int\limits_{-\infty}^{+\infty}d\nu\,\dfrac{\Gamma^{L+1}(\frac{\lambda+n+1}{2}-\frac{\beta}{2}-{\sf i}\nu)\Gamma^{L+1}(\frac{\lambda+n+1}{2}-\frac{\beta}{2}+{\sf i}\nu)}{\Gamma^{L+1}(\frac{\lambda+n+1}{2}+\frac{\beta}{2}-{\sf i}\nu)\Gamma^{L+1}(\frac{\lambda+n+1}{2}+\frac{\beta}{2}+{\sf i}\nu)}({\sf y}^{2})^{{\sf i}\nu+\beta/2}({\sf x}^{2})^{-{\sf i}\nu+\beta/2}\,$, (2.19)
which is evidently real and symmetric with respect to ${\sf x}\leftrightarrow{\sf y}$. Note that under the translation $\nu\mapsto\nu-{\sf i}\frac{\beta}{2}$ contour of integration shifts to $\operatorname{Im}\nu=\frac{\beta}{2}$. It is possible to move it back to $\operatorname{Re}\nu=0$ and obtain (2.1) if $\beta<\lambda+1$, so no poles lie in the band $0<\operatorname{Im}\nu<\frac{\beta}{2}$.
Remark 2. For the case of generic $\beta$ and $D$ we denote the contour of integration in (2.18) as $\operatorname{Im}\nu=0$. However, one should note that in the special cases one should understand it as a contour, which separates the two series of singularities arising from the gamma functions in the definition of $\tau_{n,\nu}(\beta;\lambda)$ (namely, one series at $\nu={\sf i}(\frac{\lambda+n+1}{2}-\beta)+{\sf i}k$ and another at $\nu=-{\sf i}\frac{\lambda+n+1}{2}-{\sf i}k$, where $k\in\mathbb{Z}_{\geq 0}$). For more details in the cases of specific $\beta$ and $D$ see section (3.3) and Appendix D.
Now we use the definition of the Gegenbauer polynomials $C_{n}^{(D/2-1)}$ [14, 15] (see also [70], eq. (3.48)) in (2.18):
@@
where $\hat{{\sf x}}\hat{{\sf y}}=\frac{({\sf x}{\sf y})}{\sqrt{{\sf x}^{2}{\sf y}^{2}}}$. One can consider the right-hand side of this formula as the correct analytical continuation of its left-hand side for any non-integer $D$. Let us introduce the notation $\lambda=\frac{D}{2}-1$ and new parametrization [58]
$\frac{{\sf y}^{2}}{{\sf x}^{2}}=z\bar{z}\ \ ;\ \ \frac{2({\sf x}{\sf y})}{{\sf x}^{2}}=z+\bar{z}\qquad\Rightarrow\qquad\hat{{\sf x}}\hat{{\sf y}}=\frac{z+\bar{z}}{2\sqrt{z\bar{z}}}\;.$ (2.21)
Note that, in terms of conformal ratios (2.5) and in view of (2.14), this parametrization corresponds to
$u=\dfrac{z\bar{z}}{(1-z)(1-\bar{z})}\;,\qquad\qquad v=\dfrac{1}{(1-z)(1-\bar{z})}\;.$ (2.22)
Using definition (2.20), formula (2.10) and parametrization (2.21) we write (2.18) as

### v2

@@ Abstract
In the paper, a family of conformal four-point ladder diagrams in arbitrary space-time dimensions is considered. We use the representation obtained via explicit calculation using the operator approach and conformal quantum mechanics to study their properties such as symmetries, loop and dimensional shift identities. In even dimensions, latter allows one to reduce the problem to the two-dimensional case, where notable factorization holds. Additionally, for a specific choice of propagator powers, we show that the representation can be written in the form of linear combinations of classical polylogarithms (with coefficients that are rational functions) and explore the structure of the resulting expressions.
1 Introduction
@@
In recent years, significant progress has been made in the computation of scattering amplitudes in various quantum field theories (see e. g. Bern et al. (1996, 2007); Elvang and Huang (2015); Bern et al. (1994, 1995); Arkani-Hamed et al. (2010)). In calculation of physical observables, e.g. in QCD, the algoritmizability and uniformity are often prioritized over the preserving specific symmetries at intermediate steps. As a result, techniques such as integration-by-parts (IBP) Chetyrkin and Tkachov (1981) or differential equation methods Henn (2013); Remiddi (1997); Lee (2015), have undergone significant development. On the other hand, in theories with enhanced symmetries, it is sometimes possible to exploit these symmetries directly in computations. A well-known example of such a theory is $\mathcal{N}=4$ SYM, especially in the planar limit. The underlying superconformal symmetry made it possible to fully constrain tree-level MHV amplitudes in a compact form Parke and Taylor (1986); Nair (1988) even in the early years of study.
@@
Remarkably, the idea of using specific symmetries in Feynman integrals was implemented even before the establishment of IBP-reduction (which is nowadays widely used as a universal tool), leading to an equally universal approach. Gegenbauer polynomial technique Chetyrkin et al. (1980); Kotikov (1996) (see also Kotikov and Teber (2019) for a modern review) deserves special mention. It utilizes the $D$-dimensional rotational symmetry, which is the signature of all Feynman integrals. Despite challenges in its algorithmic implementation, this technique provided a possibility to obtain a series of important results Celmaster and Gonsalves (1980); Terrano (1980); Lampe and Kramer (1983) and continues to reveal deep connections with other methods Derkachev et al. (2022).
@@
However, not all symmetries of the theory can be read directly from its Lagrangian formulation. A notable example is the conjectured $\text{AdS}/\text{CFT}$ correspondence, which relates planar $\mathcal{N}=4$ SYM to string theory on $\text{AdS}_{5}\times S^{5}$. This duality in turn reveals the underlying integrable structure. For instance, in the so-called fishnet theory, which can be obtained as a deformation of planar $\mathcal{N}=4$ SYM Gürdoğan and Kazakov (2016); Grabner et al. (2018); Kazakov and Olivucci (2018); Caetano et al. (2018); Kade and Staudacher (2025), underlying integrability leads to significant results in the calculation of four-point Basso-Dixon correlators Basso and Dixon (2017); Basso et al. (2021); Derkachov et al. (2019); Derkachov and Olivucci (2020, 2021b, 2021a); Derkachov et al. (2021); Duhr et al. (2023, 2024) and more general correlators Gromov et al. (2019a, 2018); Chicherin and Korchemsky (2022); Basso et al. (2019); Olivucci (2021, 2024); Aprile and Olivucci (2023). Another hidden symmetry, crucial to the structure of planar $\mathcal{N}=4$ SYM, is the dual conformal symmetry Alday and Roiban (2008); Drummond et al. (2008, 2010), which together with ordinary conformal symmetry forms a Yangian symmetry Drummond et al. (2009). This structure not only constrains tree-level amplitudes but can also be extended to the loop level, as reflected in the BDS ansatz for all-loop MHV amplitudes Bern et al. (2005). Note that dual conformal symmetry also finds its application beyond the four-dimensional theories Kazakov (2014); Bork et al. (2020); Bhattacharya and Lipstein (2016); Bork et al. (2015).
@@
From a technical point of view, dual conformal symmetry manifests itself as conformal invariance of the integrals associated with dual Feynman graphs. These integrals have attracted attention since the early days of multiloop calculations and continue to be actively studied Gorishnii and Isaev (1985); Drummond et al. (2007). Conformal symmetry at the level of perturbative integrals makes it possible to use powerful analytical identities, such as e.g. the star–triangle relation Zamolodchikov (1980); Vasiliev et al. (1981); Chicherin et al. (2013) or the theory of graphical functions Brown and Schnetz (2012); Schnetz (2014), to simplify calculations dramatically. The star–triangle relation (interpreted as the Yang-Baxter equation; for a brief review see the subsection 5.3 in Isaev (2025)) allows one to consider not only individual diagrams but entire families, such as fishnet diagrams Zamolodchikov (1980) or conformal ladder diagrams in arbitrary dimensions $D$ Isaev (2003, 2008). Recall that ladder diagrams were first studied by Ussyukina and Davydychev, who derived all-loop results in four dimensions Usyukina and Davydychev (1993a, b). Later, a generalization of conformal ladder integrals were analyzed, and their analytic properties, such as single-valuedness, were explored in Drummond (2013).
@@
Note that in the paper Chicherin et al. (2013) we found the general $D$-dimensional conformal invariant solution to the Yang-Baxter equation that generalizes the $1$-dimensional solution of Isaev (2008) and the $2$-dimensional solution of Derkachov et al. (2001) and underlies Lipatov’s integrable models describing high-energy behaviour of QCD as well as integrable structures of scattering amplitudes in $\mathcal{N}=4$ SUSY Lipatov (2004, 2009); Faddeev and Korchemsky (1995). This Yang-Baxter solution was used, for example, in the investigations of the Yangian symmetries of perturbative integral in fish-net type, loom and checkerboard CFT Kazakov et al. (2025); Kazakov and Olivucci (2023); Alfimov et al. (2025).
@@
In one of our previous works Derkachov et al. (2023), we used the graph-building operator technique Gürdoğan and Kazakov (2016); Grabner et al. (2018); Kazakov and Olivucci (2018); Derkachov et al. (2019); Gromov et al. (2019b); Derkachov and Olivucci (2020, 2021b); Derkachov et al. (2022) together with the connection to conformal quantum mechanics Isaev (2003, 2008) to obtain an all-loop result for the conformal ladder and zig-zag four-point correlators in arbitrary dimensions. While our expression for ladder diagrams was fully analytical and valid in any dimension, it was formulated in terms of Gegenbauer polynomials, so despite its generality, the form of the answer may be of little practical use. Recently, remarkable progress has been made in two dimensions, where conformal ladder integrals were shown to be related to twisted partition functions Petkou (2021); Karydas et al. (2024, 2025), allowing them to be rewritten in terms of classical polylogarithms. In the present work, we show that in arbitrary even dimensions and a specific choice of propagator indices our previous representation can be systematically expressed using classical polylogarithms and rational functions. Furthermore, we verify that our representation satisfies a loop and dimensional shift identity studied in Loebbert and Stawinski (2024); Karydas et al. (2024, 2025). The latter, among other things, allows one to express the answer for $D$-dimensional diagram in terms of two-dimensional, where notable factorization holds (see Derkachov et al. (2019); Derkachev et al. (2022) for examples of such a factorization). We believe that these results can be useful not only from a practical point of view but also in revealing possible underlying symmetries, such as antipodal self-duality Dixon and Duhr (2025) in the fishnet theory which also holds for the one-loop ladder diagrams in $D=4$. Before closing this section, we would like to mention that despite the rich diversity of results the studies of conformal integrals continue to attract interest Alkalaev and Mandrygin (2025a, b); He and Jiang (2025).
@@
The paper is organized as follows. In Section 2 we review the calculation of a general family of conformal four-point ladder diagrams using the graph-building operator and conformal quantum mechanics. After discussing the most general choice of parameters for ladder diagrams, we restrict ourselves to the case described by a single parameter $\beta$. We show that the obtained representation admits a dimensional shift $D\mapsto D+2$ for a general choice of $\beta$. Additionally, we show that the combination of dimensional shift operator with the graph-building operator allows us to construct an operator that increases the number of loops $L\mapsto L+1$. In Section 3 we move to the case of ladder diagrams with the specific parameter $\beta=1$ and present the derivation of the answer containing only classical polylogarithms and rational functions. Then, we study a representation based on a two-dimensional factorized form in the case $\beta=2,3,\ldots$. The last section contains our conclusions. The appendices are devoted to a discussion of technical details: in Appendix A, we derive the answer for two-dimensional conformal ladder integrals using the representation with Gegenbauer polynomials, in Appendix B, we give an explicit check of the loop shift identity for four-dimensional diagram with $\beta=1$, Appendix C shows the properties of the answer in the arbitrary dimensional case with $\beta=1$, and in Appendix D, we discuss the details of two-dimensional factorization.
2 Conformal four-point ladder integrals in diverse dimensions
2.1 Formulation of the problem and introductory remarks
@@
To fix notation, in this section we recall some known facts (see e.g. Refs. Isaev (2003); Derkachov et al. (2023)). For massless $\varphi^{3}$ theory, we consider D-dimensional $L$-loop ladder integrals with arbitrary indices on the lines $\alpha_{k},\beta_{k},\gamma_{k}$:
@@
These integrals correspond to the momentum-space Feynman diagrams depicted in Fig. 1 (here the integrations are performed over the loop momenta $p_{i}$ (i=1,,Li=1,...,L)). The diagram given in Fig. 1 is presented in the dual form in Fig. 2. Here the integrations are performed over the boldface vertices which are placed in the boxes of the diagram in Fig. 1.
@@
Figure 2: Dual graphical representation of the integral (2.1) written in the form (2.2). The integrations are performed over the boldface vertices and each line with index $\alpha$: $x_{i}\stackrel{{\scriptstyle\alpha}}{{-\!\!\!-\!\!\!-\!\!\!-}}x_{j}$, corresponds to the propagator $x_{ij}^{-2\alpha}$
@@
For further applications, we fix the indices in the integral (2.2), which is depicted in Fig.2, by the conformal conditions for all boldface vertices Isaev (2003); Derkachov et al. (2023)
@@
When $\alpha_{k},\beta_{k},\gamma_{k}$ are positive real numbers, the condition (2.3) ensures the convergence of the integral (2.2). Under the conformal transformations $y_{\mu}\to\frac{y_{\mu}}{y^{2}}$ (inversion of the vector $y\in\mathbb{R}^{D}$; $\mu=1,...,D$ is the space index) we have for vectors $x_{i},x_{j}\in\mathbb{R}^{D}$:
@@
and in the case when the conditions (2.3) are fulfilled, the integral (2.2) transforms as follows:
@@
Thus, the function***We use the notation $(x_{ij})^{A}=(x_{ij}^{2})^{A/2}$.
@@
and $f_{(0)}=1$. Further, for simplicity we fix the parameters $\beta_{k}=D/2-\gamma_{k}\equiv\beta$ (which gives $\alpha_{k}=\beta$). In this case, formula (2.6) is simplified to
@@
and these functions (associated with the diagrams in Fig 2) give contributions to the $4$-point amplitude in general $D$-dimensional bi-scalar fishnet CFT proposed in Kazakov and Olivucci (2018, 2023). For this choice of the parameters, we found in Isaev (2003, 2008); Derkachov et al. (2023) that the generating function of the $L$-loop ladder integrals (2.2) (with special normalization (2.7)) is represented in the form of a Green’s function for conformal quantum mechanicsHere to avoid confusion with the standard notation of cross-ratios, we use the notation $\sf x$} and $\sf y$} instead of $u$ and $w$ in Derkachov et al. (2023) and $u$ and $v$ in Isaev (2003).
$\begin{array}[]{c}\displaystyle\frac{a(\beta)}{({\sf x}-{\sf y})^{2(D/2-\beta)}}\cdot\sum_{L=0}^{\infty}\bigl(g\,a(\beta)\bigr)^{L}\,f_{(L)}(u,v;\beta)=\bigl\langle{\sf x}\bigr|\,\frac{1}{\hat{p}^{2\beta}-g\,\hat{q}^{-2\beta}}\,\bigl|{\sf y}\bigr\rangle\;,\\[5.69046pt] \hat{p}^{2}=\hat{p}_{\mu}\hat{p}_{\mu}\,,\;\;\;\hat{q}^{2}=\hat{q}_{\mu}\hat{q}_{\mu}\,,\;\;\;[\hat{q}_{\mu},\;\hat{p}_{\nu}]=i\delta_{\mu\nu}\,,\;\;\;\;{\sf x}:=\frac{1}{x_{12}}-\frac{1}{x_{42}}\,,\;\;\;\;{\sf y}:=\frac{1}{x_{32}}-\frac{1}{x_{42}}\,,\\[8.5359pt] \dfrac{a(\beta)}{(x-y)^{2(D/2-\beta)}}=\langle x|\dfrac{1}{\hat{p}^{2\beta}}|y\rangle\,\quad(\forall x,y\in\mathbb{R}^{D}),\;\;\;\;\;\;a(\beta):=\dfrac{\Gamma(D/2-\beta)}{2^{2\beta}\,\pi^{D/2}\Gamma(\beta)}\;,\end{array}$ (2.8)
where $\{\hat{q}_{\mu},\;\hat{p}_{\nu}\}$ are the generators of $D$-dimensional Heisenberg algebra, and Derkachov et al. (2023)
@@
Here ${\sf x}^{\mu_{1}...\mu_{n}}$ is the traceless symmetric tensor with components that are homogeneous in ${\sf x}^{\mu}$ polynomials (see Appendix A in Derkachov et al. (2023)); the functions
@@
were introduced in Derkachov et al. (2023) as weights in the completeness condition of the eigenfunctions of operators $H_{\beta}=\hat{p}^{2\beta}\hat{q}^{2\beta}$; the operators $H_{\beta}$ form a commutative set for all $\beta$ Isaev (2003) while $\tau_{n,\nu}(\beta;\lambda)$ are eigenvalues of $H_{\beta}$ (for further details see subsection 2.3):
@@
Note also that we have relations between the variables ${\sf x},{\sf y}$ (introduced in (2.8)) and cross-ratios (2.5)
@@
The representation (2.8) of the generating function of $L$-loop ladder integrals as a Green’s function for conformal quantum mechanics is very useful since we, for example, immediately deduce the symmetry properties of $\Phi^{(\beta)}_{L}({\sf x},{\sf y})$ Isaev (2003):
@@
which also follows from (2.9). Comparing (2.9) and (2.15) we obtain Derkachov et al. (2023)
@@
Remark 1. Under the translation of the variable $\nu\mapsto\nu-{\sf i}\frac{\beta}{2}$, the function (2.18) transforms to the expression
$\Phi^{(\beta)}_{L}({\sf x},{\sf y})=$ $\,(L!\,4^{L})4^{-\beta(L+1)}\sum\limits_{n=0}^{\infty}\mu(n)\,\frac{{\sf x}^{\mu_{1}...\mu_{n}}\,{\sf y}^{\mu_{1}...\mu_{n}}}{({\sf x}^{2}\,{\sf y}^{2})^{(D/4+n/2)}}\quad\times$
$\quad\times\int\limits_{-\infty}^{+\infty}d\nu\,\dfrac{\Gamma^{L+1}(\frac{\lambda+n+1}{2}-\frac{\beta}{2}-{\sf i}\nu)\Gamma^{L+1}(\frac{\lambda+n+1}{2}-\frac{\beta}{2}+{\sf i}\nu)}{\Gamma^{L+1}(\frac{\lambda+n+1}{2}+\frac{\beta}{2}-{\sf i}\nu)\Gamma^{L+1}(\frac{\lambda+n+1}{2}+\frac{\beta}{2}+{\sf i}\nu)}({\sf y}^{2})^{{\sf i}\nu+\beta/2}({\sf x}^{2})^{-{\sf i}\nu+\beta/2}\,$, (2.19)
which is evidently real and symmetric with respect to ${\sf x}\leftrightarrow{\sf y}$. Note that under the translation $\nu\mapsto\nu-{\sf i}\frac{\beta}{2}$ the contour of integration shifts to $\operatorname{Im}\nu=\frac{\beta}{2}$. It is possible to move it back to $\operatorname{Re}\nu=0$ and obtain (2.1) if $\beta<\lambda+1$, so no poles lie in the band $0<\operatorname{Im}\nu<\frac{\beta}{2}$.
Remark 2. For the case of generic $\beta$ and $D$ we denote the contour of integration in (2.18) as $\operatorname{Im}\nu=0$. However, one should note that in the special cases one should understand it as a contour, which separates the two series of singularities arising from the gamma functions in the definition of $\tau_{n,\nu}(\beta;\lambda)$ (namely, one series at $\nu={\sf i}(\frac{\lambda+n+1}{2}-\beta)+{\sf i}k$ and the other at $\nu=-{\sf i}\frac{\lambda+n+1}{2}-{\sf i}k$, where $k\in\mathbb{Z}_{\geq 0}$). For more details in the cases of specific $\beta$ and $D$, see section (3.3) and Appendix D.
Now we use the definition of the Gegenbauer polynomials $C_{n}^{(D/2-1)}$ Kotikov (1996); Kotikov and Teber (2019) (see also Derkachov et al. (2023), eq. (3.48)) in (2.18):
@@
where $\hat{{\sf x}}\hat{{\sf y}}=\frac{({\sf x}{\sf y})}{\sqrt{{\sf x}^{2}{\sf y}^{2}}}$. One can consider the right-hand side of this formula as a correct analytical continuation of its left-hand side for any non-integer $D$. Let us introduce the notation $\lambda=\frac{D}{2}-1$ and new parametrization Isaev (2003)
$\frac{{\sf y}^{2}}{{\sf x}^{2}}=z\bar{z}\ \ ;\ \ \frac{2({\sf x}{\sf y})}{{\sf x}^{2}}=z+\bar{z}\quad\Rightarrow\quad\hat{{\sf x}}\hat{{\sf y}}=\frac{z+\bar{z}}{2\sqrt{z\bar{z}}}\;.$ (2.21)
Note that in terms of conformal ratios (2.5) and in view of (2.14), this parametrization corresponds to
$u=\dfrac{z\bar{z}}{(1-z)(1-\bar{z})}\;,\quad\quad v=\dfrac{1}{(1-z)(1-\bar{z})}\;.$ (2.22)
Using definition (2.20), formula (2.10) and parametrization (2.21), we write (2.18) as